P7-2A. Variable and Absorption Costing Summarized data for 2016 (the first year of operations) for

Gorman Products, Inc., are as follows:

Sales (75,000 units) ....................................... $3,000,000

Production costs (80,000 units):

Direct material.......................................... 880,000

Direct labor ............................................ 720,000

Manufacturing overhead:

Variable ............................................... 544,000

Fixed ................................................. 320,000

Operating expenses:

Variable ............................................... 168,000

Fixed ................................................. 240,000

Depreciation on equipment ................................. 60,000

Real estate taxes ......................................... 18,000

Personal property taxes (on inventory and equipment) ............ 28,800

Personnel department expenses . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30,000

Required

a. Prepare an income statement based on full absorption costing.

b. Prepare an income statement based on variable costing.

c. Assume that you must decide quickly whether to accept a special one-time order for 1,000 units

for $30 per unit. Which income statement presents the most relevant data? Determine the appar-

ent profit or loss on the special order based solely on these data.

d. If the ending inventory is destroyed by fire, which costing approach would you use as a basis for

filing an insurance claim for the fire loss? Why?
Answer #1

ANSWER:

a:

An Income statement based on full absorption costing.

  Absorption Costing Income Statement

For the year Ended 2016

sales (a) $3,000,000
Cost of goods sold:
Beginning inventory(b) $0
Cost of goods manufactured(c)
(720,000+544,000+320,000) 1094000
Ending inventory (d) 68375
Cost of goods sold (e = b+c-d) 1,025,625
Gross margin (f = a - e) 1,974,375
operating expenses (g)
(168,000 + 240,000 ) 408,000
Income from operations (f-g)
( 1,94,375 - 408,000) 1,566,375

Compute the ending inventory value under absorption costing as shown below:

Ending inventory= Ending inventory(in jobs) x Manufacturing cost per job

= (Total jobs - Number of jobs sold) x Cost of goods manufactured / Total jobs

= (80,000-75000) x 1094000 / 80000

= 5000 x 13.675

= 68,375

Thus, the income from operation for the year 2016 based on full absorption costing is $ 1,566,375.

b)

Prepare an income statement based on varable costing as shown below:

Variable costing Income Statement

For the year ended 2016

Sales(a) $ 3,000,000
Variable cost of goods Sold:
Beginning inventory(b) $0
Variable cost of goods manufacture (c)
(720,000+544,000) 1,264,000
Ending inventory (d) 79,000
Variable cost of goods sold (e= b +c -d) 1185,000
Manufacturing margin (f = a - e) 1,815,000
Variable operating expenses (g) 168,000
Contribution margin ( h = f - g) 1,647,000

Fixed costs:

Fixed manufacturing over hears(i)   

Total fixed cost ( k =i + j)

(320,000 +2,40,000) 5,60,000

Income form operation ( h-k) 1,087,000

Compute the ending inventory value under variable costing as shown below:

Ending inventory = Ending inventory x Variable manufacturing cost per job

= (80000 - 75000) x 1,264,000 / 80000

= 5000 x 15.8 = 79000

Thus the income from operations for the year 2016 based on variable costing is $ 1,087,000

c)

For the special one time order analysis, variable income statement is useful because for the special order jobs only variable cost are incurred.

Thus, the varibale costing income statement presents the most relevant data for special one time order analysis

Special order

Sales (1000 x 30 per unit) 30,000

Variable cost of goods manufactured (b)

[(720,000+544,000) / 5000 x (1000)] 252800

Manufacturing margin ( c = a -b) 222,800

Variable operating expenses (d)

(168,000 / 75000) x 5000 11200

Net profit ( c - d) 211,600

Thus , the net profit on the special order is $ 211,600

d) If ending inventory is destroyed by fire, the full absorption method will be used because the cost per unit is higher and if closing inventory is valued by the method the value will be higher, therby getting a higher claim form the insurance company.

-------------------------------------------

If you have any query regarding the answer please as me in the comment I am here for help you. Please do not direct thumbs down just ask if you have any query. And if you like my work then please appreciates with up vote
Thank you Very much.